OpenStudy (jravenv):

There are 8 people on the debate team. In how many ways can the coach choose 4 members to send to competition? 70 1,680 40,320

OpenStudy (anonymous):

70

Mehek (mehek14):

@jobo explain

Mehek (mehek14):

You can't just give direct answers

OpenStudy (yolomcswagginsggg):

...................__ ............./´¯/'...'/´¯¯`·¸ ........../'/.../..../......./¨¯\ ........('(...´...´.... ¯~/'...') .........\.................'...../ ..........''...\.......... _.·´ ............\..............( BROFIST ...........

OpenStudy (anonymous):

\[\left(\begin{matrix}8 \\ 4\end{matrix}\right) = 70\]

jimthompson5910 (jim_thompson5910):

8 people 4 slots (A,B,C,D) 8 choices for slot A 7 for slot B 6 for slot C 5 for slot D 8*7*6*5 = 1680 choices total IF order mattered, but it doesn't since order doesn't matter, we divide by 4! = 24 to get 1680/24 = 70 the result of 70 is equal to 8 C 4 as JoBo wrote above